OpenStudy (anonymous):

simplify square root of 6 + 2(sqaure root of 6) A. 3(square root of 6) B. -(square root of 6) C. 3(square root of 12)

OpenStudy (anonymous):

do u mean \[6+2\sqrt{6}\]

OpenStudy (anonymous):

\[\sqrt{6} + 2\sqrt{6}\]

OpenStudy (cwrw238):

similar to adding x + 2x

OpenStudy (cwrw238):

treat √6 as if it was x

OpenStudy (anonymous):

\[2\sqrt{6+6} \[2\sqrt{12} \[2\sqrt{3+9} equals \[5\sqrt{3}

OpenStudy (anonymous):

none of these are one of the options

OpenStudy (cwrw238):

x + 2x = 3x similarly: sqrt6 + 2 sqrt6 = 3 sqrt6 A